# Reasoning Large Language Model Errors Arise from Hallucinating Critical Problem Features

### v1

@@ 6.3 Results
Plots of error rates by type and false-uncolorable attributability for both models on 12v6c, and of edge hallucination rate for Grok 3 MB (high) thanks to its exposed CoT, can be seen in Figure 6. Overall error rates are no more than a few percent for any model-frame combination, but they are much higher than these models’ error rates on the simpler 8v4c problem set. As with models in general on 8v4c, errors are dominated by the false-uncolorable type. Rates of detected edge hallucinations for Grok 3 MB (high) are considerably higher for both frames on 12v6c than on 8v4c, and most false-uncolorable responses contain edge hallucinations that would make them correct (except for Grok 3 MB (high) Math, where the rate is 9 out of 24 responses). For all four model-frame combinations, the most frequently hallucinated edge is (10, 11), or the equivalent (Kathy, Larry) for Friends; the rest of the frequency distribution varies, but as with 8v4c, it is always distinctly non-uniform.
@@ 7 Discussion
We have shown that a variety of RLLMs from different developers and with release dates spanning most of RLLM history are prone to hallucinate edges in (sufficiently complex) graph coloring problems, which can lead them to answer the problems incorrectly. This phenomenon occurs even if the problems are not stated explicitly in terms of graph coloring, and it persists and may even grow more frequent as problem complexity increases. The phenomenon is unexpected in the sense that all and only the graph’s true edges are listed in the prompt, which should not ever leave any of the models’ context windows given the lengths of the prompts and responses. This may indicate that RLLMs have issues distinguishing appropriately between information in the prompt and the potentially less reliable information in their own in-progress chain of thought; if so, they may benefit from a feature akin to positional encoding that explicitly demarcates prompt tokens vs. chain-of-thought tokens, if they do not already have it. The asymmetry in which edges they tend to hallucinate may also reflect weaknesses in how they represent numbers or orderings, with higher-indexed entities becoming more easily conflated with one another.
@@ 7.1 Limitations
We do not test all publicly available RLLMs, and we only test one problem type (graph coloring) as a representative of the broader category of logic tasks. However, the existence and behavior of edge hallucinations are consistent enough in the RLLMs we do test that we expect them to occur in any other RLLMs designed similarly to them, and we find it likely that the design flaws that enable edge hallucination in graph coloring also affect at least some other problem types on which the RLLMs have received comparably thorough or less thorough training.
@@ A.1 Purpose
In their graph coloring experiments, Heyman and Zylberberg [2025] categorize colorable problems using a numerical estimate of how easy they are to successfully color “greedily" one vertex at a time without ever backtracking on coloring choices, naming this measure the “greedy score". They find that among colorable problems, both o1-mini and DeepSeek-R1’s error rates increase drastically as greedy score decreases (i.e. as greedy coloring becomes harder), more sharply than with the traditional LLMs they test. They take this as evidence that though the two RLLMs represent a large improvement in step-by-step reasoning ability, their gains in possibility space exploration have not been as great. However, the correlation they observe is computed across problems with varying edge counts, and high edge count also correlates positively with error among colorable problems. We want to see if low greedy score remains related to high error when edge count is controlled for.
@@ A.2 Methods
We construct a validation problem set by randomly sampling 1000 8-vertex graphs with 19 edges and 1000 with 20 edges, for a total of 2000, for use in 4-coloring problems. We then calculate the greedy score $g$ for all of the colorable graphs and divide them into the same three categories Heyman and Zylberberg [2025] use: $g\geq 0.9$, $0.90.5\leq g<$ 0.9, and $g<0.5$, containing 250, 676, and 211 problems respectively. We randomly sample 250 of the 676 problems from the middle category to even out the category sizes, leaving a total of 711 problems (19 edges: 145, 154, and 112 for the three categories; 20 edges: 105, 96, and 99). We prompt each of o1-mini and DeepSeek-R1 with each problem five times each, using the same methodology as in Section 2, and only in the Math frame.
@@ A.3 Results
Plots of error rates for each combination of model, edge count, and greedy score category can be seen in Figure 7. Contrary to the expectations of Heyman and Zylberberg [2025], for every combination of model and edge count, error rate tends to be lower on problems with lower greedy score, which should be harder to find valid colorings for if the models’ primary area of weakness was systematic possibility space exploration in the form of backtracking or other non-greedy strategies. Less surprisingly, however, error rate is always at least slightly higher for 20-edge problems than 19-edge problems when model and greedy score category are held constant. We do not know why greedy score and error rate distinctly correlate in the opposite-of-expected direction, rather than not distinctly correlating at all. We also cannot guarantee that this pattern holds for other models, frames, or edge/vertex/color counts with a diversity of greedy scores, but neither do we have any particular reason to expect this experimental setup to be unrepresentative.

### v3

@@ 6.3 Results
Plots of error rates by type and false-uncolorable attributability for both models on 12v6c, and of edge hallucination rate for Grok 3 MB (high) thanks to its exposed CoT, can be seen in Figure 6. Overall error rates are no more than a few percent for any model-frame combination, but they are much higher than these models’ error rates on the simpler 8v4c problem set. As with models in general on 8v4c, errors are dominated by the false-uncolorable type. Rates of detected edge hallucinations for Grok 3 MB (high) are considerably higher for both frames on 12v6c than on 8v4c, and most false-uncolorable responses contain edge hallucinations that would make them correct (except for Grok 3 MB (high) Math, where the rate is 9 out of 24 responses). For all four model-frame combinations, the most frequently hallucinated edge is (10, 11), or the equivalent (Kathy, Larry) for Friends; the rest of the frequency distribution varies, but as with 8v4c, the distribution is always distinctly non-uniform.
@@ 7 Discussion
We have shown that a variety of RLLMs from different developers and with release dates spanning the majority of RLLM history are prone to hallucinate edges in (sufficiently complex) graph coloring problems, which can lead them to answer the problems incorrectly. This phenomenon occurs even if the problems are not stated explicitly in terms of graph coloring, and it persists and appears to even grow more frequent as problem complexity increases. The presence of such a straightforward input-conflicting hallucination phenomenon in LLMs specialized for solving reasoning problems is striking, and given the lengths of the prompts and responses it cannot be explained away by context window limitations. Instead, it may indicate that RLLMs have issues distinguishing appropriately between information in the prompt and the potentially less reliable information in their own in-progress chain of thought; if so, they may benefit from a feature akin to positional encoding that explicitly demarcates prompt tokens vs. chain-of-thought tokens, if they do not already have it. The asymmetry in which edges they tend to hallucinate may also reflect weaknesses in how they represent numbers or orderings, with higher-indexed entities becoming more easily conflated with one another.
We also know that the hallucination phenomenon we identify is not exclusive to graph coloring, owing to the supplementary experiments we describe in Appendix D. We generate stable matching problems in which students must be assigned to dorm rooms in groups of 3, and find that a significant source of RLLM error on them is misrepresentation of students’ ratings for each other – most often transposing A’s rating for B onto B’s rating for A, suggesting weaknesses in representing directionality.
@@ 7.1 Limitations
We do not test all publicly available RLLMs, and we only test two problem types (graph coloring and stable 3-matching) as representatives of the category of logic tasks. However, the existence and behavior of the hallucinations we observe are consistent enough in the RLLMs we do test that we expect them to occur in any other similarly designed RLLMs, and we find it unlikely that the design flaws that enable these hallucinations do not also affect at least some other problem types on which the RLLMs have received comparably thorough or less thorough training.
@@ A.1 Purpose
In Heyman and Zylberberg [2025], we categorize colorable graph coloring problems using a numerical estimate of how easy they are to successfully color “greedily" one vertex at a time without ever backtracking on coloring choices, naming this measure the “greedy score". We find that among colorable problems, both o1-mini and DeepSeek-R1’s error rates increase drastically as greedy score decreases (i.e. as greedy coloring becomes harder), more sharply than with the traditional LLMs we test. We initially took this as evidence that though the two RLLMs represent a large improvement in step-by-step reasoning ability, their gains in possibility space exploration have not been as great. However, the correlation we observed is across problems with varying edge counts, and high edge count also correlates positively with error among colorable problems. Here, we want to see if low greedy score remains related to high error when edge count is controlled for.
@@ A.2 Methods
We construct a validation problem set by randomly sampling 1,000 8-vertex graphs with 19 edges and 1,000 with 20 edges, for a total of 2,000, for use in 4-coloring problems. We then calculate the greedy score $g$ for all of the colorable graphs and divide them into the same three categories we use in Heyman and Zylberberg [2025]: $g\geq 0.9$, $0.5\leq g<0.9$, and $g<0.5$, containing 250, 676, and 211 problems respectively. We randomly sample 250 of the 676 problems from the middle category to even out the category sizes, leaving a total of 711 problems (19 edges: 145, 154, and 112 for the three categories; 20 edges: 105, 96, and 99). We prompt each of o1-mini and DeepSeek-R1 with each problem five times each, using the same methodology as in Section 2, and only in the Math frame.
@@ A.3 Results
Plots of error rates for each combination of model, edge count, and greedy score category can be seen in Figure 7. Contrary to our earlier hypothesis, for every combination of model and edge count, error rate tends to be lower on problems with lower greedy score, which should be harder to find valid colorings for if the models’ primary area of weakness was systematic possibility space exploration in the form of backtracking or other non-greedy strategies. Less surprisingly, however, error rate is always at least slightly higher for 20-edge problems than 19-edge problems when model and greedy score category are held constant. We do not know why greedy score and error rate distinctly correlate in the opposite-of-expected direction, rather than not distinctly correlating at all. We also cannot guarantee that this pattern holds for other models, frames, or edge/vertex/color counts with a diversity of greedy scores, but neither do we have any particular reason to expect this experimental setup to be unrepresentative.
@@ Appendix C Errors on 4v2c
Appendix D Rating Hallucinations in Stable 3-Matching Problems
After completing the experiments on graph coloring problems that we describe in this work’s main body, we took interest in using similar methodology to analyze RLLM behavior on a different type of reasoning problem, to test our hypothesis that phenomena akin to edge hallucination generalize beyond graph coloring and to potentially gain a broader perspective on RLLM reasoning from the effects (or lack thereof) on RLLM behavior of the differences between graph coloring and the new problem type.
The new problem type we chose is a stable matching problem in which objects must be assembled into groups of 3, rather than groups of 2 as is typical for such problems. Stable 3-matching problems have been studied in a theoretical computer science context (see e.g. Ng and Hirschberg [1991]), but to our knowledge, no previous work has discussed the specific form we use here. We frame our problem type in terms of $3n$ university students who must be grouped into $n$ dorm rooms holding 3 students each. Each student has rated each other student on a scale of 1 to 5 (note that a student can rate multiple other students equally, and two students’ ratings for each other need not be equal), and a student’s satisfaction with their room is the sum of their ratings for their two roommates. A grouping of students into rooms is unstable if and only if there exists a pair of students who could switch rooms to increase both of their satisfactions, or increase one and keep the other the same. The goal of the problem is to give an example of a stable grouping if one exists, or to report that no stable grouping exists. Because this specific problem type has not been a subject of past research, we expect RLLMs to be unfamiliar with it from their training compared to the basic form of graph coloring we have tested them on, which has been studied and discussed extensively elsewhere; we are interested in observing potential differences in RLLM behavior arising from this lower familiarity. Also of note is that we know of no algorithm for solving this problem type without fail that does not in the worst case involve checking every possible grouping, though worthwhile heuristics exist.
We generate two stable 3-matching problem sets of differing complexity and test several RLLMs with exposed chains of thought on them. We find that, similar to edge hallucination in graph coloring, the RLLMs’ chains of thought sometimes contain false claims about students’ ratings, which are most often transpositions of some student A’s rating for another student B onto B’s rating for A. These rating hallucinations can cause RLLMs to give incorrect final answers, though when the rate of incorrect final answers is high, it tends to be the case that few of them can be attributed solely to rating hallucinations, indicating major influence from other types of reasoning mistakes. These results support our hypothesis that input-conflicting hallucinations are a significant source of RLLM error across a broad range of reasoning problems, while also adding nuance to the evidence base on the role of other error sources.
D.1 Problem Sets
We procedurally generate two problem sets of 1,000 problems each: a lower-complexity one in which 6 students must be assembled into 2 groups of 3, and a higher-complexity one in which 9 students must be assembled into 3 groups of 3. We respectively call these problem sets “2g3" and “3g3" for short. To generate each problem, each student’s rating for each other student is sampled independently and uniformly at random from the integers between 1 and 5 inclusive.
In each 2g3 problem, there are $6\times(6-1)=30$ different rating values for the model to distinguish between and $\frac{6!}{(3!)^{2}\times 2!}=10$ possible groupings of students to check in the worst case. Of the 1,000 2g3 problems we generated, 288 are such that none of the 10 groupings are stable, 447 have 1 stable grouping, 210 have 2 stable groupings, and 55 have 3 or more.
In each 3g3 problem, there are $9\times(9-1)=72$ different rating values to distinguish between and $\frac{9!}{(3!)^{3}\times 3!}=280$ possible groupings to check in the worst case. Because of the large number of possible groupings, we would not expect even an “ideal" language model of a presently-realistic size to be able to solve 3g3 problems universally without fail, at least not without a means of bypassing its context window limits; however, we are still interested in studying exactly how and how much real RLLMs err on this problem set. Of the 1,000 3g3 problems we generated, 384 have 0 stable groupings out of 280, 326 have 1 stable grouping, 160 have 2, 68 have 3, and 62 have 4 or more.
D.2 Models
We test DeepSeek-R1 (through the DeepInfra online host) and Grok 3 Mini Beta (through the xAI API; both low and high reasoning settings), as well as OpenAI’s gpt-oss-120b [OpenAI, 2025a] (through DeepInfra; low, medium, and high reasoning settings) as a recently released model with good reasoning benchmark performance that exposes its chain of thought. As in the graph coloring experiments, we run all models at temperature 0, except for DeepSeek-R1, which we run at temperature 0.5 following its developers’ advice.
D.3 Prompting and Response Evaluation
We prompt each model (on each reasoning setting, where applicable) with each of our 2,000 problems once. The frame of assigning students to dorm rooms is the only frame in which we present the problems. We give the students alphabetically sequential male names (Alan, Bob, Charlie, etc.) and present their ratings for each other in a square matrix, specifying that each row shows one student’s ratings for all of the others. (See Appendix E for example prompts.) As in the graph coloring experiments, when evaluating model responses for answer correctness, we use an automatic parser and manually review and interpret any responses in which it does not detect a coherent answer.
D.4 Results: Basic Error Patterns
Figure 8: Rate of each error type for each stable matching problem set and model. Error types are grouped by whether they apply to solvable or unsolvable problems (left and right of each pair of bars, respectively), and error rates are calculated as a fraction of the trials on the corresponding problem type. Note that each subplot’s y-axis is independently scaled to its bar heights.
Analogously to our graph coloring problems, errors on our stable 3-matching problems can be divided into three different types: claiming that a problem has no stable groupings when it actually has at least one (false unsolvable), giving a grouping that is actually unstable for a problem with other groupings that are stable (wrong grouping), and giving a grouping that is actually unstable for a problem with no stable groupings (false solvable). Rates of these error types for each problem set and model are plotted in Figure 8. (The “No coherent answer" responses from gpt-oss-120b (high) on 3g3 mostly consist of cases where it abruptly stopped talking about stable matching and “forgot the question" partway through its chain of thought.) gpt-oss-120b (low) answered “no stable groupings" in about $98\%$ of its 2g3 responses and 3g3 responses each, and few of its responses check any possible groupings in detail before giving their final answers; its level of reasoning effort appears too low for this task, so we exclude it from Figure 8 and from the remainder of this analysis.
In contrast to the errors on our graph coloring problems, which are consistently dominated by the false-uncolorable type across RLLMs at the 8v4c complexity level and above, no comparably consistent pattern emerges here despite error rates on 2g3 already being higher than those on 8v4c for each model tested on both. The dominance of false-uncolorable errors in graph coloring appears to primarily result from edge hallucination, but with stable 3-matching there seems to be no similarly impactful reasoning flaw shared across models that biases their distribution of error types in a particular direction. As expected, error rates for each model are higher overall on 3g3 than on 2g3, and error rates on each problem set are lower overall for Grok 3 MB (high) and gpt-oss-120b (high) than for their respective lower-effort settings – though gpt-oss-120b’s decrease in false-unsolvable errors on 3g3 is counterbalanced by increases in the other types, and as its CoTs also become longer, it may be that the higher reasoning effort leads it to search for stable groupings for longer and thereby increases the likelihood of a false positive. Models’ error type distributions also shift when complexity increases from 2g3 to 3g3, with errors on solvable problems becoming proportionally more common, though for Grok the increase is mostly from wrong-grouping errors while for other models it is mostly from false-unsolvable errors.
D.5 Rating Hallucinations
As with the graph coloring problem type, we conducted informal analysis of a sample of erroneous responses to the stable 3-matching problems and found cases in which the model’s chain of thought made false claims about a particular student’s rating for a particular other student (e.g. saying that Fred rates Dave a 2 when he actually rates him a 3). We want to know how common this “rating hallucination" phenomenon is and to what extent it contributes to the errors in final answers that we observe. To this end, we use methodology similar to our graph coloring experiments, writing code to automatically parse rating statements in models’ stable 3-matching responses and iteratively tuning it with priority given to eliminating false positives, thus allowing the frequencies of rating hallucinations that it finds to serve as approximate lower bounds on the true frequencies.
D.6 Results: Rating Hallucinations
Figure 9: Proportional measures related to rating hallucination for each stable matching problem set and model.
Five proportions summarizing the rating hallucination results for each problem set and model can be seen in Figure 9. From left to right in each quintuplet of bars, these are: (1) the proportion of responses that have incorrect final answers (regardless of error type); (2) the proportion of responses that contain at least one false rating statement; (3) the proportion of responses with incorrect final answers that contain at least one false rating statement; (4) the proportion of those responses whose final answers would be correct if the ratings matched the falsely claimed ones; and (5) the proportion of false rating statements (with repetitions of a particular statement within a particular response excluded) that would be correct if the students were transposed; i.e. A’s rating for B is not actually $x$, but B’s rating for A is $x$. We noticed the high frequency of these transposed ratings informally and chose to measure the frequency across all responses, and it appears that in almost all conditions the vast majority of false ratings are transposed ratings (with the exception of Grok 3 MB (high) on 2g3, for which the proportion of transposed ratings is still nearly $50\%$).
Leaving proportions of transposed ratings aside, the ten combinations of problem set and model appear to fall on an approximate spectrum between two extremes. One extreme is characterized by a low rate of incorrect final answers, a low rating hallucination rate, a greatly elevated rate of rating hallucinations among incorrect responses (providing at least correlational evidence by itself that rating hallucinations are contributing to error), and a high would-be-correct rate; examples include Grok 3 MB (high) on 2g3 and gpt-oss-120b (high) on 2g3. Grok 3 MB (low) on 2g3 and gpt-oss-120b (medium) on 2g3 exhibit a weaker version of this pattern: their incorrect final answer rates are higher but still below average, their rating hallucination rates are significantly but not enormously elevated among incorrect responses, and their would-be-correct rates are below $50\%$ but still above average. The other extreme is characterized by a high incorrect final answer rate, a high rating hallucination rate, an about-equally-high rating hallucination rate among incorrect responses, and a low would-be-correct rate; examples include DeepSeek-R1 on 3g3 and Grok 3 MB (low) on 3g3. In general, models on 3g3 fall toward this end of the spectrum.
These results match our graph coloring results in that overall rates of the hallucination phenomenon are higher at the higher complexity level; we also observe a pattern across the two problem types that within a given complexity level, models with lower error rates tend to have a higher proportion of their errors attributable solely to hallucinations, as if the hallucinations are more robust to conventional strategies for improving RLLM reasoning than other types of reasoning mistakes are. However, the high-error end of the spectrum of stable 3-matching conditions is unlike anything we see in our graph coloring results; the closest point of comparison might be Claude 3.7 Sonnet and Grok 3 MB (low) on 8v4c, with high rates of false-uncolorable error but low rates of error attributability to edge hallucination compared to other models. In the high-error stable 3-matching conditions (partially akin to those graph coloring ones), we think the most likely explanation for our results is that the models frequently make rating hallucinations but also frequently make other types of reasoning mistakes – perhaps because they are less familiar with the procedure of solving this unusual problem type compared to graph coloring – which is why most responses with rating hallucinations and incorrect final answers would still be incorrect if the ratings matched the hallucinations. However, at sufficiently low problem complexity and with a sufficiently capable model, the non-hallucination mistakes diminish even more than the hallucinations do, leaving the latter to dominate.
We find in our graph coloring experiments that the distribution of frequencies at which different edges get hallucinated is consistently non-uniform on both 8v4c and 12v6c for every model-frame combination with the relevant data available, with the edges involving the highest-index vertices tending to get hallucinated most. We do not observe equally striking patterns in the frequencies at which different ratings are hallucinated in the stable 3-matching responses; in aggregate, however, we do find weak-to-moderate but highly statistically significant biases ($p<0.001$ via two-sided one-sample $t$-test) toward rating hallucinations involving higher-indexed students (e.g. Fred more likely than Alan) in most conditions. The only exceptions are Grok 3 MB (high) on both 2g3 and 3g3, for which the student indices involved in rating hallucinations were slightly low on average, though neither of these biases was statistically significant ($p>0.05$). We hypothesized that the hallucinated-edge frequency patterns in our graph coloring responses reflect weaknesses in how RLLMs represent numbers or orderings in regards to higher-indexed entities, and these results lend some support to that, though the specifics of the behavior seem to vary greatly by problem type.
Note that our findings on rating hallucination do not conflict with the previously discussed highly variable distributions of error types on these problem sets. Unlike adding an edge in a graph coloring problem, changing a rating value in a stable matching problem of this type (in a way that is not carefully targeted) does not bias the correct answer in any particular direction, since a changed rating (increased or decreased) can turn stable groupings unstable and vice versa with equal ease.
Our findings demonstrate that input-conflicting hallucinations akin to edge hallucination can contribute significantly to RLLM error on problem types other than graph coloring. The frequency of specifically transposed ratings among false ratings across all of the models we test here may also reflect weaknesses in how RLLMs represent directionality or pair orderings (an issue that would presumably not come up in the coloring of undirected graphs), though data from more tasks would be necessary to generalize confidently.
Appendix E Example Prompts for Stable 3-Matching Problems
Following are two examples of the prompts presenting stable 3-matching problems that we use in the experiments we describe in Appendix D, one from each of the 2g3 and 3g3 problem sets.
E.1 2g3
Suppose there are 6 university students - Alan, Bob, Charlie, Dave, Ethan, and Fred - who must be grouped into 2 dorm rooms holding 3 students each. Suppose each student has rated how much they would want each other student as a roommate on a scale of 1 to 5, with 5 as the highest. The ratings can be found in the following matrix, with each row showing one student’s ratings for all of the others:
A B C D E F
A X 3 5 3 5 1
B 5 X 1 4 4 3
C 4 3 X 4 4 4
D 1 2 5 X 5 4
E 2 2 2 3 X 5
F 1 3 4 5 5 X
Suppose a student’s satisfaction with their assigned room is equal to the sum of their ratings for their 2 roommates. A grouping of the students into rooms is called unstable if there exists a pair of students in different rooms who could switch rooms to increase both of their satisfactions, or increase the satisfaction of one while the other’s satisfaction stays the same. Do there exist any possible groupings that are stable (i.e. not unstable)? If there do not, write "Impossible" as the final line of your response. If there do, the final lines of your response should give an example of one in a format like the following:
Alan-Bob-Charlie
Dave-Ethan-Fred
E.2 3g3
Suppose there are 9 university students - Alan, Bob, Charlie, Dave, Ethan, Fred, George, Henry, and Ian - who must be grouped into 3 dorm rooms holding 3 students each. Suppose each student has rated how much they would want each other student as a roommate on a scale of 1 to 5, with 5 as the highest. The ratings can be found in the following matrix, with each row showing one student’s ratings for all of the others:
A B C D E F G H I
A X 2 3 4 1 5 1 1 5
B 5 X 2 4 5 1 4 3 2
C 4 2 X 4 1 5 1 1 4
D 2 1 5 X 3 1 5 2 5
E 4 3 2 3 X 2 5 1 5
F 1 4 2 3 1 X 4 3 5
G 2 1 4 4 5 5 X 2 3
H 5 1 5 3 3 3 2 X 3
I 3 4 4 5 3 2 2 5 X
Suppose a student’s satisfaction with their assigned room is equal to the sum of their ratings for their 2 roommates. A grouping of the students into rooms is called unstable if there exists a pair of students in different rooms who could switch rooms to increase both of their satisfactions, or increase the satisfaction of one while the other’s satisfaction stays the same. Do there exist any possible groupings that are stable (i.e. not unstable)? If there do not, write "Impossible" as the final line of your response. If there do, the final lines of your response should give an example of one in a format like the following:
Alan-Bob-Charlie
Dave-Ethan-Fred
George-Henry-Ian